OpenStudy (anonymous):

In how many different ways can a panel of 12 jurors and 2 alternates be chosen from a group of 15 prospective jurors?

OpenStudy (anonymous):

same as the number of ways you can exclude on person from a group of 15, namely 15 ways

OpenStudy (anonymous):

*one person

OpenStudy (anonymous):

oh maybe we are counting jurors differently from alternates?

OpenStudy (anonymous):

not clear from the question we can do that too if you like

OpenStudy (anonymous):

Im pretty sure this is a combination question

OpenStudy (anonymous):

yeah i know,the question is this do we count jurors differently than alternates

OpenStudy (anonymous):

if we do, then the answer is \[\binom{12}{15}\times \binom{3}{2}\]

OpenStudy (anonymous):

the number of ways you can choose 12 from a set of 15, times the number of ways you can choose 2 from a set of 3

OpenStudy (anonymous):

yeah we have to count them both

OpenStudy (anonymous):

ok do you know how to compute those numbers?

OpenStudy (anonymous):

\(\binom{12}{15}\) is sometimes written as \(_{15}C_{12}\)

OpenStudy (anonymous):

i have even seen \(^{15}C_{12}\)

OpenStudy (anonymous):

no I'm trying to figure out how to on the calulator but I havent really learned it all yet . can it be done? or is it done all by hand>

OpenStudy (anonymous):

you can do it with a calculator, but these two are easy

OpenStudy (anonymous):

\[\binom{3}{2}=3\] since there are 3 ways to pick 2 out of 3

OpenStudy (anonymous):

15nCr12 * 3nCr2

OpenStudy (anonymous):

i think its that but I dont know how to compute it

OpenStudy (anonymous):

\[\binom{15}{12}=\binom{15}{3}=\frac{15\times 14\times 13}{3\times 2}\] cancel first multiply last

OpenStudy (anonymous):

you can do it instantly with wolfram http://www.wolframalpha.com/input/?i=%2815+choose+3%29*%283+choose+2%29

OpenStudy (anonymous):

what's a juror?

OpenStudy (anonymous):

so the one thousand is the answer?? that seems so low

OpenStudy (anonymous):

ahhh I see!!!!!

OpenStudy (anonymous):

ok ok I get it